The measure of the acute angles of a right triangle are jn the ratio 5:7. Find the measure of the smallest angle.
Answer by amarjeeth123(570)

Let the acute angles of the triangle be 5x and 7x respectively.
5x+7x+90=180
12x=90
x=90/12.
The smallest angle is 5*90/12=450/12=37.5 degrees.